# Accretion onto supermassive and intermediate mass black holes in cosmological simulations

### v1

@@ 2.3 A new model for gas accretion onto MBHs
In this work we explore a novel way to calculate the accretion rate onto massive black holes while keeping all other aspects of the simulation the same. We note that our motivation for introducing this model is threefold:
@@
Establish a better connection between cosmological simulations and simulations of smaller-scale nuclear gas inflows. To this end, we define the region of surrounding gas from which the inflow is calculated with respect to a fixed proper distance from the MBH 444As opposed to a (time) variable distance determined by a fixed weighted number of enclosed neighbouring cells commonly used in galaxy formation simulations., and allow for a radius-dependent mass inflow rate.
@@
Absorb physically plausible scalings in model parameters that can be varied and constrained when confronted with data while keeping the approach as simple as possible. We notably do not aim for a method that closely approximates a specific physical mechanism driving inflows, since we deem it unlikely such a model would be appropriate in all cases in large cosmological simulations that follow halos of a wide range of masses across cosmic time.
@@
Develop a model that is ready to be extended and used with future modelling techniques. This includes a model for unresolved accretion disks, algorithmic changes to the time-integration more in line with the employed fluid-dynamics solver and compatibility with multi-fluid modelling approaches (Weinberger & Hernquist 2023).
@@ 2.3.1 Fixed inflow scale
Unlike the SPH-like kernel estimate used in the IllustrisTNG model, we define the surroundings of the black hole as a sphere with fixed proper radius irrespective of the number of cells in this sphere. We use a radius of pc for L12.5n768 and kpc for L50n768, and factors of and larger values for the n384 and n192 simulations, respectively. This is several times the smoothing length of collisionless particles at high redshift, and only slightly smaller than the smoothing length at (see Table 1). We use a top-hat kernel over all cells in a sphere with radius to estimate physical quantities. If there are no cells in this sphere, we increase to include at least one. This happens only for of black holes at high redshift. Due to a fixed target mass of cells, these are the cases where the gas density and thus the accretion rate is low and not significant for the overall growth. Starting at redshift , needs to be increased more frequently, reaching up to several tens of percent at , likely due to the evacuation of the central region by kinetic AGN feedback. An illustration of the different kernels is shown in Figure 1.
@@ 2.3.2 Timescale of gas inflow
The main idea of our accretion estimate is to divide the available gas mass within a fixed aperture by the characteristic timescale of accretion from these scales. Which process sets this timescale at galactic scales is an open question. Different processes might be relevant in different regimes: while Hopkins & Quataert (2010) investigate a post-merger scenario and argue that the transport of angular momentum via torques is the driving factor, Gaspari et al. (2013) investigate an efficiently cooling halo centre and argue that gas fuelling is dominated by cold clouds that chaotically fall in and lose their angular momentum via collisions; i.e., hydrodynamically. Studies of AGN feedback regulated isolated galaxy clusters show that in this regime, the halo cooling rate is limiting the accretion (e.g. Meece et al. 2017; Ehlert et al. 2023). Recent work by Hopkins et al. (2024a) indicates that magnetic fields are a crucial aspect for accretion in quasar environments on sub-pc scales (however, neglecting the effect of AGN quasar feedback in their simulation).
For our model, we focus on the regimes most relevant for mass growth; i.e., gas-rich environments with short cooling times. The Bondi accretion formula describes a spherically symmetric non-radiative accretion flow (Bondi 1952). This is inadequate especially in a regime where radiative cooling is dominating the dynamics (Gaspari et al. 2013), as well as in regimes where the interstellar medium treatment in the form of an effective equation of state interacts with the black hole accretion by providing an artificial lower limit to the sound speed and upper limit to the density at fixed pressure. Previous works have pointed this out and applied corrections in terms of a density dependent boost factor (Booth & Schaye 2009). In this work, we drop the use of the Bondi formula entirely and start from a different approach. We assume that in the dominant growth phases:
@@
Radial pressure gradients counteracting the inflow of gas can be neglected. In the case of a star-forming, multi-phase interstellar medium with the different phases in hydrostatic equilibrium (as assumed for dense, star-forming gas in the simulation; Springel & Hernquist 2003), this is a reasonable assumption for the colder phases that dominate the mass budget.
@@
Angular momentum loss is efficient, leading to an inflow rate scaling with the dynamical, or free-fall time (Hopkins & Quataert 2011, their eq. 69),
@@
Since we intend to use our accretion estimate in large volume, and consequently low resolution simulations, we do not attempt to estimate or its scaling, but, as a starting point, treat it as a constant. We consequently assume the accretion timescale to be proportional to the free-fall time onto the black hole with mass ,
similar to what has been used in previous works on galaxy clusters (Li & Bryan 2014; Ehlert et al. 2023) and in cosmological zoom simulations of galaxies (Wellons et al. 2023). Note that we only use the MBH mass, not the enclosed mass of gas and stellar component in this timescale for simplicity and numerical robustness555Since the stellar mass scales with black hole mass, its contribution can be to a degree absorbed in the normalization factor later on. In the case of an unresolved accretion disk model, we use the combined mass of black hole and accretion disk..
@@ 2.3.3 Inflow rate estimate and its time integration
We change the time integration from a single estimate of the total inflow rate per MBH to an integration on every gas cell which is subsequently added up to the black hole accretion rate. This allows us to treat the inflow as a source term in the Euler equation and integrate it on its local timestep. In case of equal timesteps of all cells in the surroundings of a MBH, both integration methods are equivalent. The inflowing gas mass onto a MBH, , is thus
@@
Here, and are the radial inflow scaling and the normalization, respectively, and are treated as free parameters. This is slightly different from Guo et al. (2023) who suggest a scaling with Bondi-radius that would replace the speed of light by the sound speed at infinity . In either case, the scaling with a characteristic radius introduces a dependence on the MBH mass that can be controlled by varying . Thus, in practice the model has two free parameters: one for the normalization and one for the MBH mass dependence.
In the following we present 2 inflow models to explore the impact of different mass dependences, one without inflow scaling as the simplest case, and one with the radial scaling of (Guo et al. 2023), . In order to find the normalization, different values for in dex steps were tested and the most similar black hole mass density evolution at or slightly below the tng model chosen. We leave a detailed calibration to future work. The employed model parameters are listed in Table 2.
@@ 2.3.4 Accretion disk model
In two of the 4 new models presented in this work, we additionally use a model for unresolved accretion disks adopted from Wellons et al. (2023). To do this, we add the inflowing mass to an accretion disk mass , which itself fuels the accretion onto the black hole,
Note that is the dynamical time of the accretion disk, which is different from the free-fall time from resolved scales. This model represents the Shakura & Sunyaev (1973) disk model, however, is not intended to precisely reflect the complex accretion disk physics, but rather to broadly explore the impact of the use or the absence of an accretion disk model might have on the cosmic MBH population. While the most obvious will be a time-averaging, with fluctuations in the inflow rate on timescales below being smoothed over in the accretion rate . Possible secondary effects on feedback remain less clear. It furthermore raises conceptual questions when applying an Eddington limit to the accretion rate, which we do throughout this work.
@@ 2.3.5 Eddington limited accretion
One notable consequence of introducing an accretion disk model in global simulations is the question of where the Eddington limit has to be applied. Since the luminosity originates from the inner accretion disk and is proportional to the innermost accretion rate, it makes sense to limit this rate; i.e. the mass inflow rate in the unresolved accretion disk . The inflow rate estimate discussed earlier, however, is no longer directly linked to the resulting luminosity. This implies that there is no motivation to limit this inflow rate, , and we consequently refrain from doing so if the accretion disk model is active. A more detailed model might, however, account for the emitted radiation and its effect on the inflow rate, effectively describing unresolved feedback moderated accretion rates. In the spirit of keeping the model simple, and due to the need for a good theoretical understanding of how emitted radiation might modulate the inflow rate, we leave such developments for future work. In the simulations without an accretion disk model, we limit the inflow rate by rescaling , i.e., the contribution of every cell to a given black hole inflow rate, by a constant factor should the inflow rate estimate exceed the Eddington limit.
@@ 3 Results
In the following we compare the new accretion prescription with the existing IllustrisTNG model (label tng). In this new model, we use a fixed aperture to determine the gas mass and the resulting free-fall time using only the black hole mass, normalized with a constant assuming a constant mass inflow rate with (label ff). As an alternative prescription, we use the same fixed aperture, but a radially dependent inflow rate scaling with and normalization (label mod ff). For both these models, we run simulations with an additional accretion disk model to assess its impact (labels ff disk and mod ff disk, respectively). In the following, we study the impact of these models onto the SMBH population in the L50n768 simulations.
3.1 Supermassive black hole evolution over cosmic time
@@
One key difference of the different inflow rate estimates is their scaling with black hole mass. Since the black hole mass function builds up over time, we would therefore expect this mass dependence to translate into a different redshift dependence between the models. Figure 2 shows the cosmic black hole accretion rate density (BHARD) as a function of redshift for the different models. The tng model in grey has the steepest rise at high redshift, but also drops off by about an order of magnitude toward low redshift. The fixed free-fall efficiency model (ff), on the other hand, leads to a shallower rise and relatively small decline of the BHARD at low redshift. We note, however, that the low-redshift decline is steeper for runs of the same model using a higher normalization (not shown here). The mod ff model sits in between the tng and ff models, both in terms of the mass dependence of the accretion rate and in the shape of the corresponding BHARD.
@@
The bottom panel of Fig. 2 shows the corresponding black hole mass density (BHMD) as a function of redshift. The final black hole mass densities differ only by a factor of about 2 (after calibrating to it), while first differences emerge at . At higher redshift, the cosmic mass densities are almost identical, likely facilitated by the use of the same MBH seed model, but also due to very similar BHARDs above redshift . For reference, we show the observationally inferred black hole accretion rate densities integrated up from the Buchner et al. (2015) luminosity functions, assuming a constant radiative efficiency of as well as the mass densities form Marconi et al. (2004); Shankar et al. (2004). While the redshift evolution is in broad agreement with the tng and mod ff simulations, it is important to note that assumptions of radiative efficiency can substantially shift the data.
@@
Fig. 3 shows the distributions of MBH masses, AGN bolometric luminosities, and Eddington ratios (i.e. the accretion rate relative to the Eddington limit, which is the maximum black holes can accrete in our model). The different rows indicate the different redshifts. The mass functions (left column) show an earlier buildup of the high mass end for the tng model at and a flattening between and that develops later for the ff model. While the mass function matches the fit presented in Shankar et al. (2004) fairly well due to our calibration to the black hole mass density, the mass functions show an increased number of massive black hole compared to Schulze et al. (2015) for the tng and mod ff models. For the bolometric luminosity (middle column) and Eddington ratio (right column) distributions, properties that depend on the instantaneous accretion rate, the changes in the distributions due to a different inflow rate estimate are remarkably small. Hardly any differences arise due to the introduction of the accretion disk model. Overall, the change with redshift is more significant than the difference between models at a given redshift. For reference, we compare the luminosity functions with fits from Shen et al. (2020, their model A). We note, however, that the quasar luminosity function is mostly fit to quasars more rare and more luminous than hosted in our L50 box.
@@
Figure 4, left column, shows the fraction of luminous AGN (X-ray luminosity ) vs. host galaxy stellar mass. There is a general expectation for more massive galaxies to host higher luminosity AGNs, thus the fraction of active galaxies increasing with stellar mass. This is indeed the case with the transition happening between and at and with only minor differences between the accretion models. For , the transition is more gradual, with a substantial fraction of higher mass galaxies not hosting a luminous AGN. The TNG model shows a characteristic plateau in AGN fraction around M⊙, similar to previously reported analysis of the IllustrisTNG simulations (Habouzit et al. 2019). Interestingly, the accretion model variations do exhibit a different behaviour with the mod ff model showing a drop at stellar mass of around M⊙ and ff a substantially weaker suppression. Compared to observations of Gallo et al. (2010); Miller et al. (2015)666Data retrieved from the analysis script of Tremmel et al. (2024)., all our models overpredict AGN in galaxies of stellar mass around M⊙, likely due to overly optimistic seeding in the IllustrisTNG model.
@@
To fully understand this behaviour, it is useful to consider the transition of feedback modes that is kept identical in all models: there is a mass-dependent transition from thermal (high Eddington ratio) to kinetic (low Eddington ratio) feedback mode (Weinberger et al. 2017). Since the kinetic mode is the by far more impactful feedback mode (Weinberger et al. 2018) it reduces not only the star formation rate of the host galaxy but also the accretion rate of the MBH. Consequently the bolometric luminosity of these AGN drops, in many cases below our threshold of an active galaxy, thus leading to a drop at precisely the location where most galaxies have their most massive MBH in kinetic mode (right panel of Figure 4). This effect is also key for the galaxy population (see Appendix A).
@@
As the ff model has a shallower scaling of the accretion rate with black hole mass compared to tng, we expect that, keeping the overall growth and environment the same, low-mass black holes grow more rapidly while massive black holes grow slower. Assuming an unaltered stellar mass buildup, this would lead to a shallower slope in MBH-galaxy scaling relations. Fig. 5 gives the black hole mass - stellar mass (left) and black hole mass - velocity dispersion (right) relations for the different models, showing median, 10th and 90th percentiles. The dashed line indicates the Greene et al. (2020) fit to MBHs at (their fit excluding upper limits). In the relations, the slope of ff is indeed shallower with a power-law slope below unity, while the tng and mod ff models show a close to unity slope above a certain mass range, but still slightly shallower than Greene et al. (2020). At higher redshift, tng and mod ff show relations below the versions for intermediate masses, while they approach their values at the highest masses. The ff model shows consistently lower mass ratios at higher redshifts. None of the models presented here show overmassive black hole mass - stellar mass relations at high redshift (Pacucci et al. 2023).
3.2 Intermediate mass black hole accretion at high redshift

### v2

@@ 2.3 A new model for gas accretion onto MBHs
In this work, we explore a novel way to calculate the accretion rate onto MBHs, while keeping all other aspects of the simulation the same. We note that our motivation for introducing this model is threefold. We aim to:
@@
establish a better connection between cosmological simulations and simulations of smaller-scale nuclear gas inflows. To this end, we defined the region of surrounding gas from which the inflow is calculated with respect to a fixed proper distance from the MBH 555As opposed to a (time) variable distance determined by a fixed weighted number of enclosed neighbouring cells commonly used in galaxy formation simulations., allowing for a radius-dependent mass inflow rate.
@@
absorb physically plausible scalings in model parameters that can be varied and constrained when confronted with data, while keeping the approach as simple as possible. Notably, we did not aim to obtain a method that closely approximates a specific physical mechanism driving inflows. This is because we deemed it unlikely that such a model would be appropriate in all cases in such large-scale cosmological simulations that follow the halos of a wide range of masses across cosmic time.
@@
develop a model that is ready to be extended and used with future modelling techniques. This includes a model for unresolved accretion disks, algorithmic changes to the time-integration more in line with the employed fluid-dynamics solver, and compatibility with multi-fluid modelling approaches (Weinberger & Hernquist 2023).
@@ 2.3.1 Fixed inflow scale
Unlike the SPH-like kernel estimate used in the IllustrisTNG model, we define the surroundings of the BH as a sphere with fixed proper radius irrespective of the number of cells in this sphere. We used a radius of pc for L12.5n768 and kpc for L50n768, and larger values for the n384 and n192 simulations, by a factor of and respectively. This is several times the smoothing length of collisionless particles at high redshift and only slightly smaller than the smoothing length at (see Table 1). We used a top-hat kernel over all cells in a sphere with radius, , to estimate physical quantities. If there are no cells in this sphere, we increased to include at least one. This happens only for of BHs at high redshift. Due to a fixed target mass of cells, these are the cases where the gas density and, thus, the accretion rate is low and not significant for the overall growth. Starting at redshift , needs to be increased more frequently, reaching up to several tens of percent at , likely due to the evacuation of the central region by kinetic AGN feedback. An illustration of the different kernels is shown in Figure 1.
@@ 2.3.2 Timescale of gas inflow
The primary aim of our accretion estimate is to divide the available gas mass within a fixed aperture by the characteristic timescale of accretion from these scales. The question of which specific process sets this timescale at galactic scales remains open. Different processes might be relevant in different regimes; while Hopkins & Quataert (2010) investigated a post-merger scenario and argue that the transport of angular momentum via torques is the driving factor, Gaspari et al. (2013) investigated an efficiently cooling halo centre and argue that gas fuelling is dominated by cold clouds that chaotically fall in and lose their angular momentum via collisions (i.e. hydrodynamically). Studies of AGN feedback-regulated isolated galaxy clusters show that in this regime, the halo cooling rate is limiting the accretion (e.g. Meece et al. 2017; Ehlert et al. 2023). A recent work by Hopkins et al. (2024a) indicated that magnetic fields are crucial for accretion in quasar environments on sub-pc scales (neglecting, however, the effect of AGN quasar feedback in the simulation).
For our model, we focus on the regimes most relevant for mass growth, namely, gas-rich environments with short cooling times. The Bondi accretion formula describes a spherically symmetric non-radiative accretion flow (Bondi 1952). This is inadequate especially in a regime where radiative cooling is dominating the dynamics (Gaspari et al. 2013), as well as in regimes where the ISM treatment in the form of an effective equation of state interacts with the BH accretion by providing an artificial lower limit to the sound speed and upper limit to the density at fixed pressure. Previous works have pointed this out and applied corrections in terms of a density dependent boost factor (Booth & Schaye 2009). In this work, we dropped the use of the Bondi formula entirely and start from a different approach. We assumed that in the dominant growth phases:
@@
radial pressure gradients counteracting the inflow of gas can be neglected. In the case of a star-forming, multi-phase ISM with the different phases in hydrostatic equilibrium (as assumed for dense, star-forming gas in the simulation; Springel & Hernquist 2003), this is a reasonable assumption for the colder phases that dominate the mass budget.
@@
angular momentum loss is efficient, leading to an inflow rate, , scaling with the dynamical or free-fall time, (Hopkins & Quataert 2011, their Eq. 69), as per
@@
Since we intend to use our accretion estimate in the context of large-volume samples and, consequently, in low-resolution simulations, we did not attempt to estimate or its scaling. Instead, we treated it as a constant. We consequently assume the accretion timescale to be proportional to the free-fall time onto the BH with mass, ,
This is similar to what was used in previous works on galaxy clusters (Li & Bryan 2014; Ehlert et al. 2023) and in cosmological zoom simulations of galaxies (Wellons et al. 2023). We note that we only use the MBH mass, not the enclosed mass of gas and stellar component in this timescale for simplicity and numerical robustness666Since the stellar mass scales with BH mass, its contribution can be to a degree absorbed in the normalisation factor later on. In the case of an unresolved accretion disk model, we used the combined mass of BH and accretion disk..
Furthermore, we did not explicitly model any mass outflow rates driven by unresolved stellar (Hopkins et al. 2024a; Partmann et al. 2025) or AGN feedback (Ostriker et al. 2010; Choi et al. 2012; Cochrane et al. 2023; Farcy et al. 2025). While the normalisation factor absorbs effects linear to the enclosed mass, non-linear effects are not taken into account.
@@ 2.3.3 Inflow rate estimate and its time integration
We changed the time integration from a single estimate of the total inflow rate per MBH to an integration on every gas cell which is subsequently added up to the BH accretion rate. This allowed us to treat the inflow as a source term in the Euler equation and integrate it on its local time step. In the case where there are equal time steps for all cells in the surroundings of a MBH, both integration methods are equivalent. The inflowing gas mass onto a MBH, , is thus
@@
Here, and are the radial inflow scaling and the normalisation, respectively, and are treated as free parameters. This is slightly different from Guo et al. (2023) who suggest a scaling with Bondi-radius that would replace the speed of light by the sound speed at infinity . In either case, the scaling with a characteristic radius introduces a dependence on the MBH mass that can be controlled by varying . Thus, in practice the model has two free parameters: one for the normalisation and one for the MBH mass dependence.
In the following we present 2 inflow models to explore the impact of different mass dependences, one without inflow scaling as the simplest case, and one with the radial scaling of (Guo et al. 2023), . In order to find the normalisation, different values for in dex steps were tested and the most similar result for the black hole mass density (BHMD) evolution at or slightly below the tng model chosen. We leave a detailed calibration to future work. The employed model parameters are listed in Table 2.
@@ 2.3.4 Accretion disk model
In two of the four new models presented in this work, we additionally use a model for unresolved accretion disks adopted from Wellons et al. (2023). To do this, we add the inflowing mass to an accretion disk mass , which itself fuels the accretion onto the BH,
We note that is the dynamical time of the accretion disk, which is different from the free-fall time from resolved scales. This model represents the Shakura & Sunyaev (1973) disk model, however, is not intended to precisely reflect the complex accretion disk physics, but rather to broadly explore the impact of the use or the absence of an accretion disk model might have on the cosmic MBH population. While the most obvious will be a time-averaging, with fluctuations in the inflow rate on timescales below being smoothed over in the accretion rate . Possible secondary effects on feedback remain less clear. It furthermore raises conceptual questions when applying an Eddington limit to the accretion rate, as we have done throughout this work.
@@ 2.3.5 Eddington limited accretion
One notable consequence of introducing an accretion disk model in global simulations is the question of where the Eddington limit has to be applied. Since the luminosity originates from the inner accretion disk and is proportional to the innermost accretion rate, it makes sense to limit this rate; i.e. the mass inflow rate in the unresolved accretion disk . The inflow rate estimate discussed earlier, however, is no longer directly linked to the resulting luminosity. This implies that there is no motivation to limit this inflow rate, , and we consequently refrain from doing so if the accretion disk model is active. A more detailed model might, however, account for the emitted radiation and its effect on the inflow rate, effectively describing unresolved feedback moderated accretion rates. In the spirit of keeping the model simple, and due to the need for a good theoretical understanding of how emitted radiation might modulate the inflow rate, we leave such developments for future work. In the simulations without an accretion disk model, we limit the inflow rate by rescaling , namely, rescaling the contribution of every cell to a given BH inflow rate by a constant factor in cases where the inflow rate estimate exceed the Eddington limit.
@@ 3 Results
In the following, we compare the new accretion prescription with the existing IllustrisTNG model (label tng). In this new model, we use a fixed aperture to determine the gas mass and the resulting free-fall time using only the BH mass, normalised with a constant assuming a constant mass inflow rate with (label ff). As an alternative prescription, we use the same fixed aperture, but a radially dependent inflow rate scaling with and normalisation (label mod ff). For both these models, we run simulations with an additional accretion disk model to assess its impact (labels ff disk and mod ff disk, respectively). In the following, we study the impact of these models onto the SMBH population in the L50n768 simulations.
3.1 SMBH evolution over cosmic time
@@
One key difference of the different inflow rate estimates is their scaling with BH mass. Since the BH mass function builds up over time, we would therefore expect this mass dependence to translate into a different redshift dependence between the models. Figure 2 shows the cosmic BHARD as a function of redshift for the different models. The tng model in grey has the steepest rise at high redshift, but also drops off by about an order of magnitude toward low redshift. The fixed free-fall efficiency model (ff), on the other hand, leads to a shallower rise and relatively small decline of the BHARD at low redshift. We note, however, that the low-redshift decline is steeper for runs of the same model using a higher normalisation, (not shown here). The mod ff model sits in between the tng and ff models, both in terms of the mass dependence of the accretion rate and in the shape of the corresponding BHARD.
@@
The bottom panel of Fig. 2 shows the corresponding BHMD as a function of redshift. The final BHMD results differ only by a factor of about 2 (after calibrating to it), while first differences emerge at . At higher redshift, the cosmic mass densities are almost identical, likely facilitated by the use of the same MBH seed model, but also due to very similar BHARDs above redshift . For reference, we show the observationally inferred BHARD integrated up from the Buchner et al. (2015) luminosity functions, assuming a constant radiative efficiency of as well as the mass densities form Marconi et al. (2004); Shankar et al. (2004). While the redshift evolution is in broad agreement with the tng and mod ff simulations, it is important to note that assumptions of radiative efficiency can substantially shift the data.
@@
Fig. 3 shows the distributions of MBH masses, AGN bolometric luminosities, and Eddington ratios (i.e. the accretion rate relative to the Eddington limit, which is the maximum BHs can accrete in our model). The different rows indicate the different redshifts. The mass functions (left column) show an earlier buildup of the high mass end for the tng model at and a flattening between and that develops later for the ff model. While the mass function matches the fit presented in Shankar et al. (2004) fairly well due to our calibration to the BHMD, the mass functions show an increased number of MBHs compared to Schulze et al. (2015) for the tng and mod ff models. For the bolometric luminosity (middle column) and Eddington ratio (right column) distributions, properties that depend on the instantaneous accretion rate, the changes in the distributions due to a different inflow rate estimate are remarkably small. Hardly any differences arise due to the introduction of the accretion disk model. Overall, the change with redshift is more significant than the difference between models at a given redshift. For reference, we compare the luminosity functions with fits from Shen et al. (2020, their model A). We note, however, that the quasar luminosity function is mostly fit to quasars more rare and more luminous than hosted in our L50 box.
@@
Figure 4, left column, shows the fraction of luminous AGN (X-ray luminosity ) versus host galaxy stellar mass. There is a general expectation for more massive galaxies to host higher luminosity AGNs, thus the fraction of active galaxies increasing with stellar mass. This is indeed the case with the transition happening between and at and with only minor differences between the accretion models. For , the transition is more gradual, with a substantial fraction of higher mass galaxies not hosting a luminous AGN. The TNG model shows a characteristic plateau in AGN fraction around M⊙, similar to previously reported analysis of the IllustrisTNG simulations (Habouzit et al. 2019). Interestingly, the accretion model variations do exhibit a different behaviour with the mod ff model showing a drop at stellar mass of around M⊙ and ff a substantially weaker suppression. Compared to observations of Gallo et al. (2010); Miller et al. (2015)777Data retrieved from the analysis script of Tremmel et al. (2024)., all our models overpredict AGN in galaxies of stellar mass around M⊙, likely due to overly optimistic seeding in the IllustrisTNG model.
@@
To fully understand this behaviour, it is useful to consider the transition of feedback modes that is kept identical in all models: there is a MBH mass-dependent transition from thermal (high Eddington ratio) to kinetic (low Eddington ratio) feedback mode (Weinberger et al. 2017), which implicitly translates to a relatively sharp transition in stellar mass. Since the kinetic mode is the by far more impactful feedback mode (Weinberger et al. 2018) it reduces not only the star formation rate of the host galaxy but also the accretion rate of the MBH. Consequently the bolometric luminosity of these AGN drops, in many cases below our threshold of an active galaxy, thus leading to a drop at precisely the location where most galaxies have their most massive MBH in kinetic mode (right panel of Figure 4). This effect is also key for the galaxy population (see Appendix A).
@@
As the ff model has a shallower scaling of the accretion rate with BH mass compared to tng, we expect that, keeping the overall growth and environment the same, low-mass BHs grow more rapidly, while massive BHs grow at a slower pace. Assuming an unaltered stellar mass buildup, this would lead to a shallower slope in MBH-galaxy scaling relations. Fig. 5 gives the BH mass-stellar mass (left) and BH mass-velocity dispersion (right) relations for the different models, showing median, 10th and 90th percentiles. The dashed line indicates the Greene et al. (2020) fit to MBHs at (their fit excluding upper limits). In the relations, the slope of ff is indeed shallower with a power-law slope below unity, while the tng and mod ff models show a close to unity slope above a certain mass range, but still slightly shallower than Greene et al. (2020). At higher redshift, tng and mod ff show relations below the versions for intermediate masses, while they approach their values at the highest masses. The ff model shows consistently lower mass ratios at higher redshifts. None of the models presented here show over-massive BH mass-stellar mass relations at high redshift (Pacucci et al. 2023).
3.2 IMBH accretion at high redshift